# Tool Unlearning for Tool-Augmented LLMs

### v1

@@ Abstract
Tool-augmented large language models (LLMs) are often trained on datasets of query-response pairs, which embed the ability to use tools or APIs directly into the parametric knowledge of LLMs. Tool-augmented LLMs need the ability to forget learned tools due to security vulnerabilities, privacy regulations, or tool deprecations. However, “tool unlearning” has not been investigated in unlearning literature. We introduce this novel task, which requires addressing distinct challenges compared to traditional unlearning: knowledge removal rather than forgetting individual samples, the high cost of optimizing LLMs, and the need for principled evaluation metrics. To bridge these gaps, we propose ToolDelete, the first approach for unlearning tools from tool-augmented LLMs. It implements three key properties to address the above challenges for effective tool unlearning and introduces a new membership inference attack (MIA) model for effective evaluation. Extensive experiments on multiple tool learning datasets and tool-augmented LLMs show that ToolDelete effectively unlearns randomly selected tools, while preserving the LLM’s knowledge on non-deleted tools and maintaining performance on general tasks.
@@ 1 Introduction
Tool-augmented Large Language Models (LLMs) can use external tools such as calculators (Schick et al., 2023), Python interpretors (Gao et al., 2023), APIs (Tang et al., 2023), or AI models (Patil et al., 2023) to complement the parametric knowledge of vanilla LLMs and enable them to solve more complex tasks (Schick et al., 2023; Patil et al., 2023). They are often trained on query-response pairs, which embed the ability to use tools directly into parameters.
Despite the growing adoption of tool-augmented LLMs, the ability to selectively unlearn tools has not been investigated. In real-world applications, tool unlearning is essential for addressing critical concerns such as security, privacy, and model reliability. For example, consider a tool-augmented LLM deployed in a healthcare system and trained to use APIs for handling patient data. If one of the APIs is later flagged as insecure due to a vulnerability that could expose sensitive information and violate regulations like HIPAA, tool unlearning is necessary to ensure that the LLM can no longer invoke the insecure API. Similarly, when tools undergo major updates, such as the Python transformers package moving from version 3 to version 4, tool unlearning becomes essential to prevent the LLM from generating outdated or erroneous code. The goal of this work is to address this gap by investigating tool unlearning and providing a solution for this overlooked yet essential task.
@@
Removing skills requires modifying the parameters of LLMs, a process that is computationally expensive and can lead to unforeseen behaviors (Cohen et al., 2024; Gu et al., 2024). In addition, existing membership inference attack (MIA) techniques, a common evaluation method in machine unlearning to determine whether specific data samples were part of training data, are inadequate for evaluating tool unlearning, as they focus on sample-level data rather than tool-based knowledge.
To address these challenges, we propose ToolDelete, the first tool unlearning algorithm for tool-augmented LLMs, which satisfies three key properties for effective tool unlearning: tool knowledge removal, which focuses on removing any knowledge gained on tools marked for unlearning; tool knowledge retention, which focuses on preserving the knowledge gained on other remaining tools; and general capability retention, which maintains LLM’s general capability on a range of general tasks such as text and code generation using ideas from task arithmetic (Ilharco et al., 2023; Barbulescu & Triantafillou, 2024). In addition, we develop LiRA-Tool, an adaptation of the Likelihood Ratio Attack (LiRA) (Carlini et al., 2022) to tool unlearning, to assess whether tool-related knowledge has been successfully unlearned. Our contributions are:
@@
Extensive experiments on multiple datasets and tool-augmented LLMs show that ToolDelete outperforms existing general and LLM-specific unlearning algorithms by $+$ in accuracy on forget tools and retain tools. In addition, it can save 74.8% of training time compared to retraining, handle sequential unlearning requests, and retain 95+performance in low resource settings.
Figure 1: Tool Unlearning and the proposed ToolDelete approach. (a): Illustration of tool learning and tool unlearning. Learned tools may be requested to be unlearned due to many reasons, such as tools being insecure, restricted, or deprecated. (b): Differences between tool unlearning and traditional sample unlearning, in terms of objective and training data. (c): Proposed method ToolDelete. We encourage the unlearned model $f^{\prime}$ to follow the tool-free LLM $f_{0}$ which has never seen $T_{f}$ before. Meanwhile, we maintain its ability on $T_{r}$ and general tasks by matching the capabilities of tool-augmented model $f$ through task arithmetic.
@@ 3.1 Tool Knowledge Deletion
$\mathop{\mathbb{missing}}{E}_{t_{i}\in\mathcal{T}_{f}}[g(f_{0},t_{i})-g(f^{\prime},t_{i})]\geq 0.$ (1)
@@ 3.2 Tool Knowledge Retention
$\mathop{\mathbb{missing}}{E}_{t_{m}\in\mathcal{T}_{r}}[g(f,t_{m})-g(f^{\prime},t_{m})]=\epsilon,$ (2)
@@ 3.3 General Capability Retention via Task Arithmetic
$\mathop{\mathbb{missing}}{E}_{t_{g}\in\mathcal{T}_{G}}[g(f_{0},t_{g})-g(f^{\prime},t_{g})]=\epsilon,$ (3)
@@
We propose to use task arithmetic (Ilharco et al., 2023; Barbulescu & Triantafillou, 2024) as an efficient and effective approach to preserving the general capabilities of the unlearned model. Our objective is that $f^{\prime}$ retains as much general knowledge as $f_{0}$, the instruction tuned LLM trained from a randomly initialized model $f_{R}$. Let $\theta_{0}$ and $\theta_{R}$ denote the parameters of $f_{0}$ and $f_{R}$ respectively. The difference vector $\theta_{0}-\theta_{R}$ captures the direction of general knowledge acquisition. We apply this adjustment to $\theta^{\prime}$ (the parameters of $f^{\prime}$) to preserve its general knowledge:
@@
Task arithmetic is efficient, practical, effective for preserving general capabilities (Ilharco et al., 2023; Barbulescu & Triantafillou, 2024): Efficiency: the vector operation does not scale with dataset size, making it significantly more efficient than retraining on large datasets. Practicality: general capabilities obtained from pre-training and instruction tuning (Zhou et al., 2024) are often impractical to replicate due to the size and limited availability of data–even in some open-source LLMs (Touvron et al., 2023b), the actual pre-training data is not fully open-source. In addition, reintroducing general knowledge from alternative datasets can lead to data imbalances and distributional biases. Effectiveness: applying $\theta_{0}-\theta_{R}$ largely restores the foundational abilities of $f^{\prime}$, such as text generation and instruction-following, without requiring expensive and time-consuming retraining on large datasets.
@@ 3.4 Training Details
$\theta^{\prime*}=\arg\min_{\theta^{\prime}}\underbrace{\mathbb{E}_{t_{i}\in\mathcal{T}_{f}}[g(f_{0},t_{i})-g(f^{\prime},t_{i})]}_{\text{knowledge deletion of }\mathcal{T}_{f}}+\underbrace{\mathbb{E}_{t_{m}\in\mathcal{T}_{r}}[g(f,t_{m})-g(f^{\prime},t_{m})]}_{\text{knowledge retention of }\mathcal{T}_{r}},$ (5)
@@ 3.5 LiRA-Tool for Tool Unlearning Evaluation
A key challenge in evaluating tool unlearning is the lack of membership inference attack (MIA) models to determine whether a tool has been truly unlearned. Existing MIA models typically evaluate individual training samples by analyzing model loss, which is insufficient for tool unlearning. Unlike sample-level unlearning, tool unlearning focuses on removing abstract parametric knowledge of tools in $\mathcal{T}_{f}$, not just forgetting specific training samples. The key limitation of sample-based MIA is that the prompt-response pairs $(\mathcal{Q}_{f},\mathcal{Y}_{f})$ in the training set may not fully represent all aspects of a tool’s functionality. As a result, sample-level MIA may “overfit” to a limited subset of tool related prompts and fail to holistically assess whether the tool-usage capability have been fully removed from the model’s parametric knowledge.
@@
where $\mathbb{P}_{t_{i}}$ represents the shadow distribution for generating tool-learning samples for tool $t_{i}$. In practice, we use GPT-4 to generate diverse shadow samples by prompting it with various distinct instructions to ensure that the evaluation set captures more comprehensive aspects of tool knowledge than the training set. Appendix E provides more details.
@@
The key novelty in LiRA-Tool in the sue of “shadow samples,” which introduce diversity across multiple dimensions. By moving beyond limited training prompts, LiRA-Tool ensures that the model loss reflect overall tool-using ability, rather than just sample-level memorization. Our loss-ratio formulation shares similarities to previous MIAs for sample-level unlearning, such as probability distribution comparison prior- and post-unlearning (Cheng et al., 2023; Cheng & Amiri, 2025) and other adaptations of LiRA using shadow models (Kurmanji et al., 2023; Pawelczyk et al., 2024). However, to the best of our knowledge, this work is the first adaptation of LiRA for detecting tool presence in tool-augmented LLMs.
@@
Table 1: Tool unlearning performances when deleting 20% of tools on ToolAlpaca. Best and second-best performances are bold and underlined respectively. Original is provided for reference only. Results on other LLMs are shown in Appendix Table 6-7.
@@ 4 Experimental Setup
As there are no prior works on tool unlearning, we adapt the following unlearning methods to tool unlearning setting (see Appendix B for descriptions of the baselines): general unlearning approaches, including GradAscent (Golatkar et al., 2020; Yao et al., 2024), RandLabel (Graves et al., 2021), and SalUn (Fan et al., 2024b); and LLM-specific unlearning approaches, including ICUL (Pawelczyk et al., 2024), SGA (Jang et al., 2023; Barbulescu & Triantafillou, 2024), TAU (Barbulescu & Triantafillou, 2024), CUT (Li et al., 2024b), NPO (Zhang et al., 2024b), and SOUL-GradDiff (Jia et al., 2024). For ICUL (Pawelczyk et al., 2024), we randomly select one example $(q_{i},y_{i})$ from $\mathcal{T}_{f}$ and corrupt the output $y_{i}$ with randomly selected tokens. Then we concatenate this corrupted sequence with other intact sequences as the in-context demonstrations. For all other baselines, we treat all data related to $\mathcal{T}_{f}$ as unlearning examples and all data related to $\mathcal{T}_{r}$ as remaining examples. Everything else remains the same for each baseline.
@@ 5 Results
Our main results in Table 1 show that, compared to Retrain, the best-performing baseline in the general unlearning methods category, ToolDelete-SFT outperforms Retrain by 0.6, 0.3, 8.0, 2.3 absolute points on $\mathcal{T}_{T}$, $\mathcal{T}_{r}$, $\mathcal{T}_{f}$, $\mathcal{T}_{G}$ respectively. ToolDelete-DPO outperforms Retrain by 1.3, 3.3, 9.8, 1.8 absolute points across the same metrics. We note that GradAscent can effectively unlearn $\mathcal{T}_{f}$, but it negatively impacts its $\mathcal{T}_{T}$ and $\mathcal{T}_{r}$ performance. Although RandLabel and SalUn outperforms GradAscent, they still fall short on $\mathcal{T}_{G}$ compared to ToolDelete.
@@
DPO outperforms SFT by 0.7, 3.0, and 1.8 on $\mathcal{T}_{T}$, $\mathcal{T}_{r}$, $\mathcal{T}_{f}$ respectively. On $\mathcal{T}_{G}$, SFT is slightly better than DPO by 0.5 points. However, DPO takes slightly longer time to train, see Figure 4. Both optimization methods achieve superior performance over existing approaches.
@@
Efficiency is a critical aspect for unlearning. As Figure 4 in Appendix D illustrates, ToolDelete is substantially more efficient than retraining a new model from scratch–saving about 74.8% of training time on average. In addition, this efficiency gain is relatively consistent as the size of $T_{f}$ increases. ToolDelete-SFT is slightly faster than ToolDelete-DPO, as the latter requires a negative sample for each of its prompts.
@@
Why is ToolDelete effectiveness?
We attribute the performance of ToolDelete to its three key properties: (a): Tool Knowledge Removal enables targeted tool unlearning without over-forgetting, unlike GradAscent and Retrain. This is achieved by prioritizing tool knowledge-free responses over tool knowledge-intense responses so that the model forgets tool functionality without excessive degradation. (b): Tool Knowledge Retention reinforces the knowledge about remaining tools. In fact, re-exposing the model to the original training data can further strengthen their representation. (c): General Capability Retention, which maintains or even improves model’s general capabilities through an efficient and effective task arithmetic operation. Therefore, precise unlearning, retention of relevant knowledge, and overall model stability are the key factors that contribute to the performance of ToolDelete.

### v2

@@ Abstract
Tool-augmented large language models (LLMs) are often trained on datasets of query-response pairs, which embed the ability to use tools or APIs directly into the parametric knowledge of LLMs. As these models are increasingly deployed in real-world applications, there is a need for them to forget specific tools–for example, due to security vulnerabilities, privacy regulations, or tool deprecation. This work presents “tool unlearning” as a novel machine unlearning task that presents distinct challenges beyond traditional sample-level unlearning: it requires removing functional knowledge rather than individual data points, managing the high cost of LLM optimization, and developing principled evaluation metrics. To address these challenges, we propose ToolDelete, the first unlearning framework designed specifically for tool-augmented LLMs. It implements three key properties for effective tool unlearning and introduces a new membership inference attack (MIA) model for effective evaluation. Extensive experiments on multiple tool learning datasets and tool-augmented LLMs show that ToolDelete effectively unlearns both randomly selected and class-specific tools, while preserving knowledge on remaining tools and maintaining performance on general tasks.
Machine Learning, ICML
@@ 1 Introduction
Tool-augmented Large Language Models (LLMs) can use external tools such as calculators (Schick et al., 2023), Python interpretors (Gao et al., 2023), APIs (Tang et al., 2023), or AI models (Patil et al., 2023) to complement the parametric knowledge of vanilla LLMs and enable them to solve more complex tasks (Schick et al., 2023; Patil et al., 2023). They are often trained on query-response pairs, which embed the ability to use tools directly into model parameters.
Despite the growing adoption of tool-augmented LLMs, the ability to selectively unlearn tools has not been investigated. In real-world applications, tool unlearning is essential for addressing critical concerns such as security, privacy, and model reliability. For example, consider a tool-augmented LLM deployed in a healthcare system and trained to use APIs for handling patient data. If one of the APIs is later flagged as insecure due to a vulnerability that could expose sensitive information and violate regulations like HIPAA, tool unlearning is necessary to ensure that the LLM can no longer invoke the insecure API. Similarly, when tools undergo major updates, such as the Python transformers package moving from version 3 to version 4, tool unlearning becomes essential to prevent the LLM from generating outdated or erroneous code. The goal of this work is to address this gap by investigating tool unlearning and providing a solution for this crucial task.
@@
Removing skills requires modifying the parameters of LLMs, a process that is computationally expensive and can lead to unforeseen behaviors (Cohen et al., 2024; Gu et al., 2024). In addition, existing membership inference attack (MIA) techniques, a common evaluation method in machine unlearning to determine whether specific data samples were part of training data, are inadequate for evaluating tool unlearning because they focus on sample-level data rather than tool-based knowledge.
To address these challenges, we propose ToolDelete, the first tool unlearning algorithm for tool-augmented LLMs, which satisfies three key properties for effective tool unlearning: tool knowledge removal, which focuses on removing any knowledge gained on tools marked for unlearning; tool knowledge retention, which focuses on preserving the knowledge gained on other remaining tools; and general capability retention, which maintains LLM’s general capability on a range of general tasks such as text and code generation using ideas from task arithmetic (Ilharco et al., 2023; Bărbulescu & Triantafillou, 2024). In addition, we develop LiRA-Tool, an adaptation of the Likelihood Ratio Attack (LiRA) (Carlini et al., 2022; Pawelczyk et al., 2024) to tool unlearning, to assess whether tool-related knowledge has been successfully unlearned. Our contributions are:
@@
Extensive experiments on multiple datasets and tool-augmented LLMs show that ToolDelete outperforms existing general and LLM-specific unlearning algorithms by $12.5$ and $9.1$ in accuracy on forget tools and retain tools respectively. In addition, it can save 74.8% of training time compared to retraining, handle sequential unlearning requests, and retain 95% performance in low resource setting.
Figure 1: Tool Unlearning and the proposed ToolDelete approach. (a): Illustration of tool learning and tool unlearning. Learned tools may be requested to be unlearned due to many reasons, such as tools being insecure, restricted, or deprecated. (b): Differences between tool unlearning and traditional sample unlearning, in terms of objective and training data. (c): Proposed method ToolDelete. We encourage the unlearned model $f^{\prime}$ to follow the tool-free LLM $f_{0}$ which has never seen $T_{f}$ before. Meanwhile, we maintain its ability on $T_{r}$ by matching the capabilities of tool-augmented model $f$ through task arithmetic.
@@ 3.1 Tool Knowledge Deletion
$\mathop{\mathbb{E}}_{t_{i}\in\mathcal{T}_{f}}[g(f_{0},t_{i})-g(f^{\prime},t_{i})]\geq 0.$ (1)
@@ 3.2 Tool Knowledge Retention
$\mathop{\mathbb{E}}_{t_{m}\in\mathcal{T}_{r}}[g(f,t_{m})-g(f^{\prime},t_{m})]=\epsilon,$ (2)
@@ 3.3 General Capability Retention via Task Arithmetic
$\mathop{\mathbb{E}}_{t_{g}\in\mathcal{T}_{G}}[g(f_{0},t_{g})-g(f^{\prime},t_{g})]=\epsilon,$ (3)
@@
We propose to use task arithmetic (Ilharco et al., 2023; Bărbulescu & Triantafillou, 2024) as an efficient and effective approach to preserving the general capabilities of the unlearned model. Our objective is that $f^{\prime}$ retains as much general knowledge as $f_{0}$, the instruction tuned LLM trained from a randomly initialized model $f_{R}$. Let $\theta_{0}$ and $\theta_{R}$ denote the parameters of $f_{0}$ and $f_{R}$ respectively. The difference vector $\theta_{0}-\theta_{R}$ captures the direction of general knowledge acquisition. We apply this adjustment to $\theta^{\prime}$ (the parameters of $f^{\prime}$) to preserve its general knowledge:
@@
Task arithmetic is efficient, practical, effective for preserving general capabilities (Ilharco et al., 2023; Bărbulescu & Triantafillou, 2024): Efficiency: the vector operation does not scale with dataset size, making it significantly more efficient than retraining on large datasets. Practicality: general capabilities obtained from pre-training and instruction tuning (Zhou et al., 2024) are often impractical to replicate due to the size and limited availability of data–even in some open-source LLMs (Touvron et al., 2023b), the actual pre-training data is not fully open-source. In addition, reintroducing general knowledge from alternative datasets can lead to data imbalances and distributional biases. Effectiveness: applying $\theta_{0}-\theta_{R}$ largely restores the foundational abilities of $f^{\prime}$, such as text generation and instruction-following, without requiring expensive and time-consuming retraining on large datasets.
@@ 3.4 Training Details
$\theta^{\prime*}=\arg\min_{\theta^{\prime}}\underbrace{\mathbb{E}_{t_{i}\in\mathcal{T}_{f}}[g(f_{0},t_{i})-g(f^{\prime},t_{i})]}_{\text{knowledge deletion of }\mathcal{T}_{f}}+\\ \underbrace{\mathbb{E}_{t_{m}\in\mathcal{T}_{r}}[g(f,t_{m})-g(f^{\prime},t_{m})]}_{\text{knowledge retention of }\mathcal{T}_{r}},$ (5)
@@ 3.5 LiRA-Tool for Tool Unlearning Evaluation
A key challenge in evaluating tool unlearning is the lack of membership inference attack (MIA) models to determine whether a tool has been truly unlearned. Existing MIA models typically evaluate individual training samples by analyzing model loss, which is insufficient for tool unlearning. Unlike sample-level unlearning, tool unlearning focuses on removing abstract parametric knowledge of tools in $\mathcal{T}_{f}$, not just forgetting specific training samples. The key limitation of sample-based MIA is that the prompt-response pairs $(\mathcal{Q}_{f},\mathcal{Y}_{f})$ in the training set may not fully represent all aspects of a tool’s functionality. As a result, sample-level MIA may “overfit” to a limited subset of tool related prompts and fail to holistically assess whether the tool-usage capability have been fully removed from the model’s parametric knowledge (Lynch et al., 2024; Łucki et al., 2025; Hu et al., 2025).
@@
where $\mathbb{P}_{t_{i}}$ represents the shadow distribution for generating tool-learning samples for tool $t_{i}$. $P_{U}(\cdot)$ indicates the distribution of unlearned tools $T_{f}$ under the unlearned model $f^{\prime}$, while $P_{T_{r}}(\cdot)$ denotes the distribution of the retain tools $T_{r}$ under the retained model $f$. In practice, we use GPT-4 to generate diverse shadow samples by prompting it with various distinct instructions to ensure that the evaluation set captures more comprehensive aspects of tool knowledge than the training set. Appendix E provides more details.
@@
The key novelty in LiRA-Tool in the sue of “shadow samples,” which introduce diversity across multiple dimensions. By moving beyond limited training prompts, LiRA-Tool ensures that the model loss reflect overall tool-using ability, rather than just sample-level memorization. Our loss-ratio formulation shares similarities to previous MIAs for sample-level unlearning, such as probability distribution comparison prior- and post-unlearning (Cheng et al., 2023; Cheng & Amiri, 2024a) and other adaptations of LiRA using shadow models (Kurmanji et al., 2023; Pawelczyk et al., 2024). However, to the best of our knowledge, this work is the first adaptation of LiRA for detecting tool presence in tool-augmented LLMs.
@@
Table 1: Tool unlearning performances when deleting 20% of tools on ToolAlpaca. Best and second-best performances are bold and underlined respectively. Original is provided for reference only. Results on other LLMs are shown in Appendix Table 5-6.
@@ 4 Experimental Setup
As there are no prior works on tool unlearning, we adapt the following unlearning methods to tool unlearning setting (see Appendix B for descriptions of the baselines): general unlearning approaches, including GradAscent (Golatkar et al., 2020; Yao et al., 2024), RandLabel (Graves et al., 2021), and SalUn (Fan et al., 2024); and LLM-specific unlearning approaches, including ICUL (Pawelczyk et al., 2024), SGA (Jang et al., 2023; Bărbulescu & Triantafillou, 2024), TAU (Bărbulescu & Triantafillou, 2024), CUT (Li et al., 2024b), NPO (Zhang et al., 2024b), and SOUL-GradDiff (Jia et al., 2024). For ICUL (Pawelczyk et al., 2024), we randomly select one example $(q_{i},y_{i})$ from $\mathcal{T}_{f}$ and corrupt the output $y_{i}$ with randomly selected tokens. Then we concatenate this corrupted sequence with other intact sequences as the in-context demonstrations. For all other baselines, we treat all data related to $\mathcal{T}_{f}$ as unlearning examples and all data related to $\mathcal{T}_{r}$ as remaining examples. Everything else remains the same for each baseline.
@@ 5 Results
Our main results in Table 1 show that ToolDelete outperforms general unlearning baselines. Compared to Retrain, the best-performing baseline, ToolDelete-SFT achieves gains of 0.6, 0.3, 8.0, 2.3 absolute points on $\mathcal{T}_{T}$, $\mathcal{T}_{r}$, $\mathcal{T}_{f}$, $\mathcal{T}_{G}$ respectively. ToolDelete-DPO shows even stronger results, outperforming Retrain by 1.3, 3.3, 9.8, 1.8 points on the same metrics. We note that GradAscent can effectively unlearn $\mathcal{T}_{f}$, but it negatively impacts its $\mathcal{T}_{T}$ and $\mathcal{T}_{r}$ performance. Although RandLabel and SalUn outperforms GradAscent, they still fall short on $\mathcal{T}_{G}$ compared to ToolDelete.
@@
DPO outperforms SFT by 0.7, 3.0, and 1.8 on $\mathcal{T}_{T}$, $\mathcal{T}_{r}$, $\mathcal{T}_{f}$ respectively. On $\mathcal{T}_{G}$, SFT is slightly better than DPO by 0.5 points. However, DPO takes slightly longer time to train, see Figure 4 in Appendix D. Both optimization methods achieve superior performance over existing approaches.
@@
Efficiency is a critical aspect for unlearning. As Figure 4 illustrates, ToolDelete is substantially more efficient than retraining a new model from scratch–saving about 74.8% of training time on average. In addition, this efficiency gain is relatively consistent as the size of $T_{f}$ increases. ToolDelete-SFT is slightly faster than ToolDelete-DPO, as the latter requires a negative sample for each of its prompts.
@@
Why is ToolDelete effective?
We attribute the performance of ToolDelete to its three key properties: (a): Tool Knowledge Removal enables targeted tool unlearning without over-forgetting, unlike GradAscent and Retrain. This is achieved by prioritizing tool knowledge-free responses over tool knowledge-intense responses so that the model forgets tool functionality without excessive degradation. This formulation imposes the right strength of forgetting over specific tools, while existing methods may over- or under-unlearn. (b): Tool Knowledge Retention reinforces the knowledge about remaining tools. In fact, re-exposing the model to the original training data can further strengthen their representation. (c): General Capability Retention, which maintains or even improves model’s general capabilities through an efficient and effective task arithmetic operation. Therefore, precise unlearning, retention of relevant knowledge, and overall model stability are the key factors that contribute to the performance of ToolDelete.